You have a client starting a new training program, so you assess her aerobic fitness using the YMCA submaximal bike test. Use the following information to answer the questions below.

Client A
Age: 30 y/o
Weight: 145 lb.
Stage 1 (Load: 0.5kp) HR: 105 bpm
Stage 2 HR: 120 bpm
Stage 3 HR: 135bpm

a. What resistance did Client A pedal at for Stage 2? What was resistance during stage 3?

b. What is her predicted VO2max in ml/kg/min?

c. What fitness category is she in?

d. What changes would you expect to see after a 6-week aerobic training program?
Answer #1

a). Stage 2 resistance =300 kgm/min (1.0kg)

Stage 3 resistance =450 kgm/min (1.5kg)

(Taken from the standard chart available online)

b). To predict the HR max, we use the formula,

220-age= 220-30= 190

From the standard chart, we get the power as given in part a.

300kgm/min and 450kgm/min is converted to watts using the conversion factor 6.12 (divide the value by 6.12). The graph of Heart rate vs power output is plotted. On extrapolating up to HRmax (190), the peak power value on the x-axis is obtained.

heart rate vs. power 150 120 105 heart rate 30 40 50 60 70 power

The formula for calculation of VO2max :

peak power* conversion factor(for power)*( conversion factor for VO2 max)/ body weight in kg + conversion constant

166.67*6.12* 1.8 /65.77 +7 = 34.91 = approximately 35

c). for a 30-year-old female

average category > VO2max = 35-38

below average category > VO2max = 31-34

She lies approximately in the average category. (Borderline value).

d). After a six-week aerobic training program, Client A's VO2 max ( maximum rate of oxygen uptake and consumption during exercise) is likely to increase due to improvement in peak output power.